Circular No. 6662

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 970508 D. A. Frail, National Radio Astronomy Observatory; S. R. Kulkarni,

California Institute of Technology; and the members of the BeppoSAX GRB Team report: “Observations with the Very Large Array of a region that includes the new x-ray source 1SAX J0653.8+7916 (IAUC 6656) and the optical transient (IAUC 6654, 6655) have resulted in the detection of a flaring radio source. No source was detected on May 9.06 UT (3.7 hr after the initial burst) and May 9.84 at a frequency of 1.43 GHz; the 3σ upper limits are 0.14 mJy and 0.27 mJy, respectively. On May 13.96, a compact source, size < 1′′, was detected with a peak flux at 8.46 GHz of 0.43 ± 0.03 mJy. On May 15.08, the source had increased to 0.61 ± 0.04 mJy at the same frequency; the flux densities were 0.10 ± 0.05 and 0.33 ± 0.03 mJy at at 1.43 and 4.86 GHz, respectively, on this day. From the observations of May 13, and assuming that the flux density is a power-law function of the frequency, we find the slope of the power law is +1.1. The best-derived position is from the 8.46-GHz observations of May 15.08 and is α = 6^{h}53^{m}49^{s}.45, δ = +79^{o}16′19′′.5 (equinox 2000.0), with an error of order 0′′.1 in both coordinates. The position of this source coincides with the position of the optical transient reported on IAUC 6655. The optical transient has been considered to be a counterpart of GRB 970508. Based on the position coincidence and the unusual flaring activity, we suggest that this is the radio counterpart of GRB 970508. In some models, the delay between the increased flaring of the radio transient with respect to the optical transient yields a valuable diagnostic. Clearly, observations at higher radio frequencies would be most valuable.”

COMET C/1996 J1 (EVANS-DRINKWATER) Z. Sekanina, Jet Propulsion Laboratory, writes: “Standard analysis

(1982, Comets, University of Arizona Press, p. 251) of the reported separation of the two nuclei (IAUC 6653) confirms that nucleus B is the primary component and suggests tentatively the second half of July 1996 as the most probable time of splitting, at a relative deceleration of ∼ 30 units of 10^{−5} the solar attraction. The estimated uncertainty is several months, and the comet might have broken up even before discovery, although there seems no hope for detecting the secondary nucleus on the available 1996 images, which range from May 10 to July 18. Predicted separation distances and position angles for companion A relative to primary B: 1997 May 12, 83′′, 253^{o}; June 1, 88′′, 258^{o}; June 21, 94′′, 262^{o}; July 11, 101′′, 266^{o}.”

1997 May 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green